OpenStudy (anonymous):

150 = 50e^ 7x (natural logs) solve for x

OpenStudy (unklerhaukus):

\[150 = 50e^ {7x}\]\[\frac{150}{50} = e^ {7x}\]\[\ln\frac{150}{50} = \ln e^ {7x}\]\[\ln\frac{150}{50} = 7x\ln e\]\[\ln\frac{150}{50} = 7x\]\[\frac 17\ln\frac{150}{50} = x\]

OpenStudy (anonymous):

so u can cancel the ln e, thanks big help

OpenStudy (lgbasallote):

i suggest simplifying 150/50 though

OpenStudy (lgbasallote):

it's cheaper with latex

OpenStudy (unklerhaukus):

yeah \(\ln e=1\) because \[\ln x=\log_e x\] and \[e^1=e\]

OpenStudy (anonymous):

thats what i did :) thanks to both of you

OpenStudy (lgbasallote):

i didnt do anything other than make a random pun. why thank me lol